OpenStudy (anonymous):

on the blueprint of a house, 48 millimeters represents 7 meters. the length of the living room is 60 millimeters on the blueprint, what is the actual length of the living room?

OpenStudy (kira_yamato):

48 mm --> 7m (i.e. 7000mm) Scale = 48mm:7m Length of living room = 60(7)/48 = 8.75 m

OpenStudy (anonymous):

8.75 m KIRA

OpenStudy (anonymous):

i cant beat you :(

OpenStudy (kira_yamato):

At least you provided a simpler method with just now question

OpenStudy (anonymous):

60/48 = 1.25 x 7 = 8.75

OpenStudy (anonymous):

Yeah I guess (: You're pretty smart.

OpenStudy (kira_yamato):

I'm not... I just write down my thoughts of the question in the answer box